# Residual-based Chebyshev filtered subspace iteration for sparse Hermitian eigenvalue problems tolerant to inexact matrix-vector products with applications to Kohn-Sham density functional theory

### v4

@@ 4 Results and Discussion
We present a detailed evaluation of the accuracy and efficiency achieved by the proposed residual-based Chebyshev filtered subspace iteration (R-ChFSI) method for solving sparse real symmetric and complex Hermitian generalized eigenvalue problems. We compare the performance of the R-ChFSI method (Algorithm 3) with that of traditional ChFSI approach (Algorithm 2), demonstrating its improved accuracy even when employing inexact matrix-vector products. Additionally, we investigate the computational performance of our proposed method on the Intel Data Centre GPU Max Series, utilising the Aurora supercomputing system, demonstrating its ability to maintain high accuracy while achieving greater computational efficiency.
@@
For all our benchmark studies reported in this work, we consider the sparse matrix eigenvalue problems that arise from higher-order finite-element (FE) discretisation of Kohn-Sham density functional theory (DFT) [chenAdaptiveFiniteElement2014, schauerAllelectronKohnSham2013, tsuchidaAdaptiveFiniteelementMethod1996, tsuchidaElectronicstructureCalculationsBased1995, dasDFTFE10Massively2022, motamarriDFTFEMassivelyParallel2020, paskFiniteElementMethods2005, motamarriHigherorderAdaptiveFiniteelement2013], widely used in ab initio modelling of materials. The Kohn-Sham DFT equations, discretised using a non-orthogonal basis, are a canonical example of the large generalised Hermitian nonlinear eigenvalue problem [^3] The generalised eigenproblem arises in the finite-element discretisation of DFT equations because of the non-orthogonality of the FE basis functions. that motivates polynomial filtered-based methods. The nonlinear eigenvalue problem is solved as a sequence of linear eigenvalue problems using a self-consistent procedure (SCF). Each iteration of SCF requires a solution of the generalised Hermitian sparse eigenproblem for the lowest $n$ eigenpairs with $n<<m$, with $n$ proportional to the number of electrons in the material system and $m$ being the number of FE basis functions. Hence, we believe that Kohn-Sham DFT provides a demanding, realistic testbed for our algorithmic claims. For the sparse generalized eigenvalue problem of the form $\textbf{A}\textbf{x}=\lambda\textbf{B}\textbf{x}$ considered here, we explore the use of diagonal approximation of B obtained by the mass lumping scheme, commonly employed in finite-element methods, to compute its inverse only during the subspace construction step within both the ChFSI and R-ChFSI methods, showing that the latter can achieve significantly lower residual tolerances.
@@
We now evaluate the accuracy and performance of the proposed R-ChFSI method for solving the aforementioned Hermitian definite generalised eigenvalue problems arising in DFT by replacing algorithm 2 in algorithm 1 by algorithm 3, We implement the proposed R-ChFSI in the open-source code DFT-FE [^4] https://github.com/dftfeDevelopers/dftfe, a massively parallel finite-element code written in C++ for first principles based materials modelling using Kohn-Sham density functional theory (DFT). To this end, we consider a set of three benchmark material systems employing periodic boundary conditions, with the sparse matrix dimensions 1.728 million, 2.567 million, 85.766 million seeking 3000, 7000 and 13500 eigenpairs, respectively. These details are summarised in table 1. The three benchmark systems are chosen to be representative of a wide range of typical finite-element discretised DFT problems, considering the following factors: (a) differing spectral gap between wanted and unwanted eigenpsectrum, (b) differing number of target eigenpairs, and (c) both Gamma-point (real symmetric) and k-point (complex Hermitian) matrices. These distinctions matter for filtering-based eigensolvers. Systems with a small or zero spectral gap between wanted and unwanted spectra require higher amplification of the desired eigenspectrum and hence higher polynomial degree to separate near-degenerate states, while systems with a larger gap allow for lower polynomial degree and converge faster. Similarly, targeting a larger number of eigenvectors increases the cost of the Rayleigh-Ritz step (as the Rayleigh-Ritz step scales cubically, compared to the quadratic scaling of the filtering step) and alters the balance of where runtime is spent. Finally, k-point sampling of the Brillouin zone in DFT produces complex Hermitian matrices and different communication/compute patterns compared to Gamma-point runs, demonstrating that both cases ensure the conclusions apply to the typical varieties of DFT calculations encountered in practice.
@@
Table 1: Dimensions of the benchmark problems considered. These correspond to 3 material systems comprising (1) $6\times 6\times 6$ supercell of Molybdenum, (2) $12\times 12\times 12$ supercell of Silicon and (3) $15\times 15\times 15$ supercell of Carbon, each with a single vacancy.
@@
System # of DoFs (m) # of wanted eigenvectors (n) Subspace dimension
(1) 1728000 3000 3600
(2) 25672375 7000 8400
(3) 85766121 13500 16800
@@
We will now demonstrate that the R-ChFSI method converges to a significantly lower residual tolerance than the ChFSI method when the diagonal approximation is employed for approximating the inverse overlap matrix during the subspace construction. To this end we will first consider the ChFSI and R-ChFSI algorithms for the benchmark systems summarized in table 1 using various Chebyshev polynomial filter degrees and compare the residual norm ($\max_{j}r_{j}^{(i)}=\max_{j}\lVert\textbf{A}\textbf{x}_{j}^{(i)}-\epsilon_{j}^{(i)}\textbf{B}\textbf{x}_{j}^{(i)}\rVert\,$) that can be acheived for both the methods. We pick a target residual norm criteria as $\max_{j}r_{j}^{(i)}<10^{-8}$, a typical tolerance employed in DFT calculations. We subsequently proceed to analyse the behaviour of the R-ChFSI method when low-precision arithmetic is employed during the subspace construction. To this end, we benchmark the robustness and performance of the R-ChFSI algorithm when employing FP32 arithmetic. To this end, in algorithm 3 we store $\textbf{R}_{X}$ and $\textbf{R}_{\textbf{Y}}$ in FP32 and the computation of $\textbf{A}\textbf{D}^{-1}\textbf{R}_{\textbf{Y}}$ is done using FP32 arithmetic. Further, we also employ TF32 tensor cores (by using oneapi::mkl::blas::compute_mode::float_to_tf32), and we report the residual norms achieved by the FP32 and TF32 variants of R-ChFSI for our benchmark systems for various values of the Chebyshev polynomial degree. We further observe that using BF16 arithmetic does not provide any noticable performance during the filtering step and instead slows down the convergence, instead we chose to perform the computation using TF32 tensor cores and the nearest neighbour MPI communication required during FE discretized matrix multi-vector products in BF16 precision, we term this method TF32B for convenience of notation. This further improves the performance of the eigensolver, as will be demonstrated subsequently. For all benchmarking studies, we consider two cases: real symmetric and complex Hermitian eigenvalue problems.

### v5

@@ 4 Results and Discussion
We present a detailed evaluation of the accuracy and efficiency achieved by the proposed residual-based Chebyshev filtered subspace iteration (R-ChFSI) method for solving real symmetric and complex Hermitian generalized eigenvalue problems. We compare the performance of the R-ChFSI method (Algorithm 3) with that of the traditional ChFSI approach (Algorithm 2), demonstrating its improved accuracy even when employing inexact matrix-vector products. Additionally, we investigate the computational performance of our proposed method on the Intel Data Center GPU Max Series, utilizing the Aurora supercomputing system, demonstrating its ability to maintain high accuracy while achieving greater computational efficiency. We first validate the theoretical claims of section 3 on small dense eigenvalue problems with precisely controlled approximation errors, and then demonstrate the practical effectiveness of R-ChFSI on large-scale sparse eigenproblems arising in Kohn-Sham DFT.
4.1 Controlled numerical experiments on dense random matrices
Before evaluating R-ChFSI on application-scale eigenproblems, we first isolate the effect of inexact matrix–vector products on the convergence of ChFSI and R-ChFSI in a fully controlled, reproducible setting. To this end we construct a family of dense eigenvalue problems of moderate size ($m=1000$) whose eigenvalue spectra, condition numbers, and approximation errors are all prescribed exactly. Since all parameters are explicitly controlled, the experiments allow a direct, quantitative verification of the convergence bounds derived in section 3, independent of any application-specific details.
4.1.1 Problem construction and experimental setup
Spectrum of A
We generate an $m\times m$ ($m=1000$) real symmetric matrix A with a prescribed eigenvalue spectrum. Towards this, a random orthogonal matrix Q is first obtained by computing the QR factorization (via Householder reflections, LAPACK routines dgeqrf/dorgqr) of an $m\times m$ matrix whose entries are drawn i.i.d. from $\mathcal{N}(0,1)$ with a fixed random seed for reproducibility. We then set $\textbf{A}=\textbf{Q}\,\mathrm{diag}(\lambda_{1},\dots,\lambda_{m})\,\textbf{Q}^{\top}$, where the $n=10$ wanted eigenvalues are uniformly spaced in the interval $[1,4]$, i.e. $\lambda_{j}=1+3(j-1)/(n-1)$ for $j=1,\dots,n$, and the remaining mn=990m-n=990 unwanted eigenvalues are $\lambda_{j}=5+0.2\,(j-n-1)$ for $j=n+1,\dots,m$. The maximal eigenvalue is therefore $\lambda_{m}\approx 202.8$. This construction yields a spectral gap of $\delta=\lambda_{n+1}-\lambda_{n}=5.0-4.0=1.0$ between the wanted and unwanted spectra. The 2-norm condition number of A is $\kappa_{2}(\textbf{A})=\lambda_{m}/\lambda_{1}\approx 203$, which is moderate and ensures that any convergence difficulties observed are attributable to the injected noise rather than ill-conditioning of the matrix itself. The Chebyshev polynomial filter of degree $p=8$ is applied with bounds $\lambda_{\min}=\lambda_{1}-0.05=0.95$, $\lambda_{\mathrm{T}}=(\lambda_{n}+\lambda_{n+1})/2=4.5$, and $\lambda_{\max}=\lambda_{m}+0.1$. In all experiments we iterate for a maximum of $100$ outer iterations without early stopping, so that the full convergence history is recorded.
Standard eigenproblem: noise model.
For the standard eigenproblem ($\textbf{B}=\textbf{I}$), the matrix is perturbed directly: we form $\tilde{\textbf{A}}=\textbf{A}+\varepsilon\,\textbf{E}$, where E is a random symmetric matrix normalized so that $\lVert\textbf{E}\rVert\,_{2}=1$ and $\varepsilon\geq 0$ controls the perturbation magnitude. The approximate inverse is set to $\textbf{D}^{-1}=\textbf{I}$ (exact). During the Chebyshev filtering step the matrix–vector products use the perturbed $\tilde{\textbf{A}}$, while the Rayleigh–Ritz projection and residual norms are evaluated with the exact A. This models scenarios in which the matrix–vector product in the polynomial filter is executed in lower precision—for example, using mixed-precision arithmetic—while the Rayleigh–Ritz step retains full-precision arithmetic. We consider four perturbation levels, $\varepsilon\in\{0,\,10^{-4},\,10^{-3},\,10^{-2}\}$.
Generalized eigenproblem: construction of B and approximate inverse.
For the generalized eigenproblem $\textbf{A}\textbf{x}=\lambda\textbf{B}\textbf{x}$, we additionally construct a symmetric positive-definite matrix B using a second random orthogonal basis $\textbf{Q}_{B}$ (same seed as Q, yielding a shared eigenbasis so that the generalized eigenvalues are exactly $\lambda_{j}^{(\mathrm{gen})}=a_{j}/b_{j}$). The eigenvalues of B are linearly spaced in the interval $[1,5]$, i.e. $b_{j}=1+4(j-1)/(m-1)$ for $j=1,\dots,m$, giving $\kappa_{2}(\textbf{B})=b_{m}/b_{1}=5$. The exact inverse is then $\textbf{B}^{-1}=\textbf{Q}_{B}\,\mathrm{diag}(1/b_{1},\dots,1/b_{m})\,\textbf{Q}_{B}^{\top}$, and the approximate inverse is formed as
$\textbf{D}^{-1}=\textbf{B}^{-1}+\zeta\,\textbf{E},\qquad\textbf{E}=\tfrac{1}{2}(\textbf{F}+\textbf{F}^{\top}),\quad F_{ij}\sim\mathcal{N}(0,1),$
where E is a random symmetric perturbation normalized so that $\lVert\textbf{E}\rVert\,_{2}=1$, and $\zeta\geq 0$ controls the perturbation magnitude. By construction, $\lVert\textbf{D}^{-1}-\textbf{B}^{-1}\rVert\,_{2}=\zeta$ exactly. This construction directly models the scenario encountered in practice when $\textbf{B}^{-1}$ is replaced by a lumped-mass or incomplete-factorization approximations. We test four levels, $\zeta\in\{0,\,10^{-4},\,10^{-3},\,10^{-2}\}$.
4.1.2 Standard eigenproblem with perturbed matrix ($\textbf{B}=\textbf{I}$)
Figure 1 shows the convergence of ChFSI and R-ChFSI for the standard eigenproblem with matrix perturbation levels $\varepsilon\in\{0,\,10^{-4},\,10^{-3},\,10^{-2}\}$. Without any perturbation ($\varepsilon=0$) both methods converge identically, as expected from their algebraic equivalence when the same matrix is used everywhere. For every nonzero perturbation level, however, the two methods diverge sharply: ChFSI stagnates at a subspace angle and residual of $\mathcal{O}(\varepsilon)$ once the Chebyshev filter can no longer distinguish the wanted subspace from the noise introduced by $\tilde{\textbf{A}}$, while R-ChFSI converges to the exact eigenvectors of the unperturbed A down to machine precision. The explanation mirrors the generalized-problem analysis of section 3: in the residual-based recurrence, the noise from $\tilde{\textbf{A}}$ enters proportional to the current residual $\lVert\textbf{R}^{(i)}\rVert\,$, which vanishes during convergence, so the filtering error decays in tandem with the residual. In the traditional recurrence, the same noise is applied to the eigenvectors themselves, whose norm remains $\mathcal{O}(1)$, producing a persistent $\mathcal{O}(\varepsilon)$ error.
Figure 1: Convergence of ChFSI and R-ChFSI for the standard eigenproblem ($\textbf{B}=\textbf{I}$, $m=1000$, $n=10$, $p=8$, $\kappa_{2}(\textbf{A})\approx 203$, spectral gap $\delta=1$) with matrix perturbation levels $\varepsilon=\lVert\tilde{\textbf{A}}-\textbf{A}\rVert\,_{2}\in\{0,10^{-4},10^{-3},10^{-2}\}$ and exact inverse $\textbf{D}^{-1}=\textbf{I}$. The Chebyshev filter uses $\tilde{\textbf{A}}$; the Rayleigh–Ritz step and residual norms use exact A. For $\varepsilon>0$, ChFSI stagnates at a subspace angle of $\mathcal{O}(\varepsilon)$, while R-ChFSI converges to machine precision.
4.1.3 Generalized eigenproblem with approximate $\textbf{B}^{-1}$
We now consider the generalized eigenproblem $\textbf{A}\textbf{x}=\lambda\textbf{B}\textbf{x}$ with a non-trivial overlap matrix B ($\kappa_{2}(\textbf{B})=5$). Because A and B share the same eigenbasis by construction, the generalized eigenvalues $\lambda_{j}^{(\mathrm{gen})}=\lambda_{j}/b_{j}$ are fully determined by the prescribed spectra of A and B, and the exact generalized eigenpairs are computed independently via LAPACK’s dsygv to serve as ground truth. The approximation quality of $\textbf{B}^{-1}$ is controlled through $\zeta=\lVert\textbf{D}^{-1}-\textbf{B}^{-1}\rVert\,_{2}$, varying from 0 (exact inverse) to $10^{-2}$ (an approximation of this magnitude might arise from a coarse lumped-mass approximation).
Figure 2 summarises the convergence behavior. Without approximation ($\zeta=0$), both methods converge identically to residuals of order $10^{-13}$. For moderate perturbation levels the two methods diverge markedly. At $\zeta=10^{-4}$ and $\zeta=10^{-3}$, ChFSI stagnates at residuals of $\mathcal{O}(\zeta)$, since the Chebyshev filtering step applies the operator $\textbf{D}^{-1}\textbf{A}$ repeatedly, amplifying the perturbation $\textbf{D}^{-1}-\textbf{B}^{-1}$ at each degree of the recurrence. R-ChFSI, on the other hand, reaches residuals of order $10^{-14}$ in both cases, confirming that the residual-based recurrence successfully decouples the converged accuracy from the quality of the approximate inverse.
For the largest perturbation tested ($\zeta=10^{-2}$), the contrast between the two methods remains striking: ChFSI stagnates at a subspace angle of $\mathcal{O}(\zeta)$, whereas R-ChFSI continues to converge monotonically, reaching a subspace angle of approximately $6\times 10^{-8}$ after $100$ iterations. Although the sufficient convergence condition eq. 5 is only marginally satisfied at this noise level, the residual-based recurrence still succeeds in driving the filtering error down in tandem with the residual. This observation highlights that R-ChFSI substantially extends the range of approximation quality $\zeta$ for which convergence is achievable, allowing machine-precision results with coarser approximations than ChFSI can tolerate.
Figure 2: Convergence of ChFSI and R-ChFSI for the generalized eigenproblem ($m=1000$, $n=10$, $p=8$, $\kappa_{2}(\textbf{B})=5$, spectral gap $\delta=1$) with varying approximation quality $\zeta=\lVert\textbf{D}^{-1}-\textbf{B}^{-1}\rVert\,_{2}\in\{0,10^{-4},10^{-3},10^{-2}\}$. For $\zeta\leq 10^{-3}$, R-ChFSI converges to machine precision while ChFSI stagnates at a residual of $\mathcal{O}(\zeta)$. Even at $\zeta=10^{-2}$, R-ChFSI continues to converge (subspace angle $\approx 6\times 10^{-8}$ at iteration $100$) while ChFSI stagnates at $\mathcal{O}(\zeta)$.
4.1.4 Verification of the convergence condition
The controlled setting of these experiments allows us to verify the convergence condition eq. 5 directly by computing both sides of the inequality at each iteration. We focus on the generalized problem with $\zeta=10^{-3}$, since this noise level is large enough to cause ChFSI to stagnate while still permitting R-ChFSI to converge, making the contrast most informative.
Figure 3 (left panel) plots the filtering error $\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,=\lVert\textbf{B}^{1/2}\boldsymbol{\Delta}_{p}^{(i)}\rVert\,$ as a function of the iteration index. For ChFSI, this quantity decreases initially (as the subspace rotates towards the eigenspace and the filter becomes more effective) but saturates at $\mathcal{O}(\zeta)$. It cannot decrease further because the filtering step accumulates a noise contribution of magnitude $\mathcal{O}(\zeta)$ at every degree of the Chebyshev recurrence, and this contribution is independent of the current subspace quality. For R-ChFSI, the filtering error tracks the residual norm and decreases monotonically to machine precision, since the recurrence filters residuals rather than eigenvectors. The effective noise injected at each step is therefore proportional to $\lVert\textbf{R}^{(i)}\rVert\,$ and vanishes as convergence progresses.
The right panel of fig. 3 displays the left- and right-hand sides of the convergence condition eq. 5. The left-hand side, $\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,$, is a constant determined solely by the spectrum and the filter parameters. For ChFSI, the right-hand side $\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,\!\left(\sec{\angle(\mathcal{S}^{(i)},\mathcal{S})}+\csc{\angle(\mathcal{S}^{(i)},\mathcal{S})}\right)$ initially decreases as the subspace rotates into alignment with the eigenspace but then rises as the subspace angle diminishes further—the trigonometric factor $\sec\theta+\csc\theta$ grows while the filtering error $\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,$ stagnates at $\mathcal{O}(\zeta)$. The two curves cross, after which the condition is violated and the subspace angle can no longer decrease—the residual stagnates. For R-ChFSI, the filtering error decreases proportionally to the residual norm, which counteracts the growth of the trigonometric factor, so the right-hand side remains below the left-hand side throughout and convergence proceeds unimpeded to machine precision. These observations confirm, in a quantitative and fully reproducible manner, the central theoretical claim of section 3: the residual-based reformulation ensures that the filtering error decays in tandem with the residual, thereby preventing the stagnation inherent in the traditional Chebyshev recurrence.
Figure 3: Diagnostic verification of the convergence condition (eq. 5) for the generalized eigenproblem with $\zeta=10^{-3}$, $m=1000$, $n=10$, $p=8$. Left panel shows the filtering error $\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,=\lVert\textbf{B}^{1/2}\boldsymbol{\Delta}_{p}^{(i)}\rVert\,$. For ChFSI it saturates at $\mathcal{O}(\zeta)$, whereas for R-ChFSI it decays monotonically to machine precision. Right panel plots the left-hand side (constant Chebyshev gap $\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,$) and the right-hand side ($\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,(\sec\theta+\csc\theta)$) of eq. 5. The condition is eventually violated for ChFSI, whereas it remains satisfied for R-ChFSI throughout.
@@ 4.1.4 Verification of the convergence condition
These controlled experiments establish that the advantages of R-ChFSI over ChFSI arise directly from the algebraic structure of the residual-based recurrence and its robustness to any mismatch between the operator used for filtering and the one used for the Rayleigh–Ritz projection. In the standard eigenproblem, when the Chebyshev filter uses an approximate $\tilde{\textbf{A}}$ while the projection uses exact A, ChFSI stagnates at a residual of $\mathcal{O}(\varepsilon)$ whereas R-ChFSI converges to machine precision, since the filtering noise enters the residual-based recurrence proportionally to the residual norm and therefore vanishes during convergence. In the generalized eigenproblem, where $\textbf{D}^{-1}\neq\textbf{B}^{-1}$, the same mechanism applies: ChFSI stagnates at $\mathcal{O}(\zeta)$ while R-ChFSI converges for all noise levels tested ($\zeta\leq 10^{-2}$), substantially extending the usable range of approximation quality. These observations are fully consistent with the theoretical predictions of section 3. Having validated the theory on problems with precisely controlled errors, we now proceed to evaluate R-ChFSI on large-scale sparse eigenproblems arising in Kohn-Sham DFT. For all our benchmark studies reported in the remainder of this section, we consider the sparse matrix eigenvalue problems that arise from higher-order finite-element (FE) discretization of Kohn-Sham density functional theory (DFT) [3, 39, 43, 42, 8, 33, 35, 32], widely used in ab initio modeling of materials. The Kohn-Sham DFT equations, discretized using a non-orthogonal basis, are a canonical example of the large generalized Hermitian nonlinear eigenvalue problem [^3] The generalized eigenproblem arises in the finite-element discretization of DFT equations because of the non-orthogonality of the FE basis functions. that motivates polynomial filtered-based methods. The nonlinear eigenvalue problem is solved as a sequence of linear eigenvalue problems using a self-consistent procedure (SCF). Each iteration of SCF requires a solution of the generalized Hermitian sparse eigenproblem for the lowest $n$ eigenpairs with $n<<m$, with $n$ proportional to the number of electrons in the material system and $m$ being the number of FE basis functions. Hence, we believe that Kohn-Sham DFT provides a demanding, realistic testbed for our algorithmic claims. For the sparse generalized eigenvalue problem of the form $\textbf{A}\textbf{x}=\lambda\textbf{B}\textbf{x}$ considered here, we explore the use of diagonal approximation of B obtained by the mass lumping scheme, commonly employed in finite-element methods, to compute its inverse only during the subspace construction step within both the ChFSI and R-ChFSI methods, showing that the latter can achieve significantly lower residual tolerances.
@@
We now evaluate the accuracy and performance of the proposed R-ChFSI method for solving the aforementioned Hermitian definite generalized eigenvalue problems arising in DFT by replacing algorithm 2 in algorithm 1 by algorithm 3. We note that the DFT-FE code already employs several mixed-precision strategies within the eigensolver—for example, mixed-precision in the Rayleigh–Ritz projection and the Gram–Schmidt orthogonalization. Our R-ChFSI modifications are complementary, as they target the filtering step (sparse matrix–multi-vector products in the Chebyshev recurrence), which is the dominant cost for the problem sizes of interest and was previously performed entirely in FP64. The mixed-precision strategies already present in DFT-FE therefore remain active and unaffected, while R-ChFSI unlocks additional savings in the computationally most expensive part of the eigensolver. We implement the proposed R-ChFSI in the open-source code DFT-FE [^4] https://github.com/dftfeDevelopers/dftfe, a massively parallel finite-element code written in C++ for first principles based materials modeling using Kohn-Sham density functional theory (DFT). To this end, we consider a set of three benchmark material systems employing periodic boundary conditions, with the sparse matrix dimensions 1.728 million, 2.567 million, 85.766 million seeking 3000, 7000 and 13500 eigenpairs, respectively. These details are summarized in table 1. The three benchmark systems are chosen to be representative of a wide range of typical finite-element discretized DFT problems, considering the following factors: (a) differing spectral gap between wanted and unwanted eigenspectrum, (b) differing number of target eigenpairs, and (c) both Gamma-point (real symmetric) and k-point (complex Hermitian) matrices. These distinctions matter for filtering-based eigensolvers. Systems with a small or zero spectral gap between wanted and unwanted spectra require higher amplification of the desired eigenspectrum and hence higher polynomial degree to separate near-degenerate states, while systems with a larger gap allow for lower polynomial degree and converge faster. Similarly, targeting a larger number of eigenvectors increases the cost of the Rayleigh-Ritz step (as the Rayleigh-Ritz step scales cubically, compared to the quadratic scaling of the filtering step) and alters the balance of where runtime is spent. Finally, k-point sampling of the Brillouin zone in DFT produces complex Hermitian matrices and different communication/compute patterns compared to Gamma-point runs, demonstrating that both cases ensure the conclusions apply to the typical varieties of DFT calculations encountered in practice.
@@
Table 1: Dimensions and key spectral properties of the benchmark problems considered. These correspond to 3 material systems comprising (1) $6\times 6\times 6$ supercell of Molybdenum, (2) $12\times 12\times 12$ supercell of Silicon and (3) $15\times 15\times 15$ supercell of Carbon, each with a single vacancy. The spectral gap is defined as $\delta=\lambda_{n+1}-\lambda_{n}$ where $\lambda_{n}$ and $\lambda_{n+1}$ are the largest wanted and smallest unwanted eigenvalues, respectively. Subspace fraction denotes $n/m$.
@@
System # of DoFs ($m$) # of wanted eigenvectors ($n$) Subspace dim. Subspace fraction ($n/m$) Spectral gap ($\delta$)
(1) 1728000 3000 3600 $1.7\times 10^{-3}$ $\sim 0$ (metallic)
(2) 25672375 7000 8400 $2.7\times 10^{-4}$ $\sim 0.03$ Ha
(3) 85766121 13500 16800 $1.6\times 10^{-4}$ $\sim 0.18$ Ha
@@
We will now demonstrate that the R-ChFSI method converges to a significantly lower residual tolerance than the ChFSI method when the diagonal approximation is employed for approximating the inverse overlap matrix during the subspace construction. To this end we will first consider the ChFSI and R-ChFSI algorithms for the benchmark systems summarized in table 1 using various Chebyshev polynomial filter degrees and compare the residual norm ($\max_{j}r_{j}^{(i)}=\max_{j}\lVert\textbf{A}\textbf{x}_{j}^{(i)}-\epsilon_{j}^{(i)}\textbf{B}\textbf{x}_{j}^{(i)}\rVert\,$) that can be achieved for both the methods. We pick a target residual norm criteria as $\max_{j}r_{j}^{(i)}<10^{-8}$, a typical tolerance employed in DFT calculations. We subsequently proceed to analyse the behavior of the R-ChFSI method when low-precision arithmetic is employed during the subspace construction. To this end, we benchmark the robustness and performance of the R-ChFSI algorithm when employing FP32 arithmetic. To this end, in algorithm 3 we store $\textbf{R}_{\textbf{X}}$ and $\textbf{R}_{\textbf{Y}}$ in FP32 and the computation of $\textbf{A}\textbf{D}^{-1}\textbf{R}_{\textbf{Y}}$ is done using FP32 arithmetic. Further, we also employ TF32 tensor cores (by using oneapi::mkl::blas::compute_mode::float_to_tf32), and we report the residual norms achieved by the FP32 and TF32 variants of R-ChFSI for our benchmark systems for various values of the Chebyshev polynomial degree. We further observe that using BF16 arithmetic does not provide any noticeable performance improvement during the filtering step and instead slows down the convergence. We therefore perform the computation using TF32 tensor cores and the nearest-neighbour MPI communication required during FE-discretized matrix multi-vector products in BF16 precision; we term this method TF32B for convenience of notation. This further improves the performance of the eigensolver, as will be demonstrated subsequently. For all benchmarking studies, we consider two cases, namely real symmetric and complex Hermitian eigenvalue problems.